[15] 9. By using the Laplace transform method solve the initial value problem Y + 2y + y = sint, y(0) = 0, y(0) = 0.
Answer #1

ana 24 $ =t 2 0-1)! L = ent En 1 (141) del from (xxx) we hawe LLY CUL 11 Flor + Iron(t) + 2007 y (H) = 2 (ett tet +601(1) =